what is product formula of integration

Grade:12

1 Answers

Sunil Kumar FP
10 years ago

if u=u(x),v=v(x), and thedifferentialsdu=u'(x)dxanddv=v'(x)dx, then integration by parts states that

\int u(x) v'(x) \, dx = u(x) v(x) - \int u'(x) v(x) \, dx

thanks and regards
sunil kr
askIItian faculty